OpenStudy (anonymous):

5(y+2)-y Struggling lol

OpenStudy (superdavesuper):

Hi, Mari, see I can help u the right way ;) 5(y+2) = 5*2 + 5*y and 5*y - y = 4y... Let us know if u need more help.

OpenStudy (anonymous):

ok